OpenStudy (anonymous):

x + 3y = 6 2x – y = 10 x + 2 x + 6 –x + 2 x + 2

OpenStudy (anonymous):

What are we doing with this?

OpenStudy (anonymous):

solving systems of equations

OpenStudy (mysesshou):

x + 3y = 6 2x – y = 10 2(-3y+6) -y = 10 -6y+12-y=10 -7y = -2 y= 2/7 then you can plug the Y into one of the equations and solve for x. your other writings below question are confusing.

OpenStudy (mysesshou):

x= -3*(2/7) + 6